OpenStudy (bossimbacon):

How many solutions does the equation have? 4x+1=5+2(2x-4) A. one solution B. infinite solutions C. no solution

OpenStudy (bossimbacon):

@immanuelv

OpenStudy (perl):

first expand the expression 4x + 1 = 5 + 4x - 4

OpenStudy (immanuelv):

it would be 4x +1 = 4x + 3 so it wont be infinite solutions

OpenStudy (bossimbacon):

a

OpenStudy (immanuelv):

both have the same slope

OpenStudy (immanuelv):